OpenStudy (anonymous):

Use the quadratic formula to solve the equation. Help Please! 2x^2-3x=-7

OpenStudy (jamesj):

First of all, what is the quadratic formula? If you had a quadratic of this form: \[ ax^2 + bx + c = 0 \] what is the formula for the solutions?

OpenStudy (anonymous):

a = 2 b = -3 and c = 7 plug these into the formula [-b +- sqrt(b^2-4ac)] /2a

OpenStudy (jamesj):

Right, \[ x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \] Now, write your equation in the standard form I did above. Then find the values of a, b, and c. Then put them into the equation.

OpenStudy (anonymous):

Ok this is my answer \[x= 3\pm \sqrt{47}/4\] is that the answer?

OpenStudy (anonymous):

= --3 +- sqrt (9 - 4*2*7) ---------------------- 4 = (-3/4) +- sqrt(-47) / 4

OpenStudy (anonymous):

so, is two solutions?

OpenStudy (anonymous):

the roots are complex numbers -(3/4) +- (sqrt47 * i )/ 4

OpenStudy (anonymous):

yes 2 solutions

OpenStudy (anonymous):

is there any other way to express that solution. When I try to plug my answer it doesn't give you the choice for i.

OpenStudy (anonymous):

sorry - i made one error its (3/4) not -(3/4)

OpenStudy (anonymous):

can you use the equation bottom so I can understand the answer. Is hard to look at it like that.

OpenStudy (anonymous):

ill try - i'm not very good at it - i'll draw it instead

OpenStudy (anonymous):